Two men swim in the river at the rate of 5 km/h.One is upstream a distance equal to the width of the river( half km) and back again.The other swims to a point straight across and back again.Assuming that the velocity of the river is 3 km/h.Find out the difference in time taken by them.

Grade:11

3 Answers

Aman gohel
6 years ago
Two men swim in the river at the rate of 5 km/h.One is upstream a distance equal to the width of the river( half km) and back again.The other swims to a point straight across and back again.Assuming that the velocity of the river is 3 km/h.Find out the difference in time taken by them.
Vikas TU
6 years ago
Dear Student,
The man going upstream voyages a separation of 1km and the other one as well. 
The resultant speed with which he the man going upstream voyages = Velocity of the man – speed of upstream = 5-3 = 2m/s 
Time taken by him=distance/speed = ½ =30min 
Alternate swims at a resultant speed of = √(25+9) = √34 (Because the even part is 3m/s and vertical is 5m/s) 
time taken by him = 60/√34 = 10.29 
Time contrast = 19.71 s.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Upendra Yadav
6 years ago
The man going upstream voyages a separation of 1km and the other one as well. 
The resultant speed with which he the man going upstream voyages = Velocity of the man – speed of upstream = 5-3 = 2m/s 
Time taken by him=distance/speed = ½ =30min 
Alternate swims at a resultant speed of = √(25+9) = √34 (Because the even part is 3m/s and vertical is 5m/s) 
time taken by him = 60/√34 = 10.29 
Time contrast = 19.71 s.
Cheers!!
Regards,
UPENDRA YADAV